OpenStudy (anonymous):

what is the difference among a^2, -a^2, (-a)^2, -(a^2)?

OpenStudy (anonymous):

a^2 = (-a)^2 -a^2 = -(a^2)

OpenStudy (anonymous):

\[a ^{2} = a*a\] \[-a ^{2} = -(a*a)\] \[(-a)^{2}=(-a)*(-a)\] The middle one would always give you a negative number, whereas the other two would always give you a positive number.

OpenStudy (anonymous):

a^2 = a^2 -a^2 = -1 x a^2 (-a)^2 = a^2 -(a^2) = -a^2

OpenStudy (anonymous):

Oops. Forgot the last one. It would essentially be the same as the middle one in my last reply.